### v1

# DWaste: Greener AI for Waste Sorting using Mobile and Edge Devices

Abstract
The rise of convenience packaging has led to generation of enormous waste, making efficient waste sorting crucial for sustainable waste management. To address this, we developed DWaste, a computer vision-powered platform designed for real-time waste sorting on resource-constrained smartphones and edge devices, including offline functionality. We benchmarked various image classification models (EfficientNetV2S/M, ResNet50/101, MobileNet) and object detection (YOLOv8n, YOLOv11n) using a subset of our own waste data set and annotated it using the custom tool Annotated Lab. We found a clear trade-off between accuracy and resource consumption: the best classifier, EfficientNetV2S, achieved high accuracy ($\approx 96\%$) but suffered from high latency ($\approx 0.22\text{s}$) and elevated carbon emissions. In contrast, lightweight object detection models delivered strong performance (up to 77% $7\%\text{ mAP}$) with ultra-fast inference ($\approx 0.03\text{s}$) and significantly smaller model sizes ($7\text{MB}$), making them ideal for real-time, low-power use. Model quantization further maximized efficiency, substantially reducing model size and VRAM usage by up to $75\%$. Our work demonstrates the successful implementation of "Greener AI" models to support real-time, sustainable waste sorting on edge devices.
Keywords: Model Quantization, Edge Computing, Object Detection, Waste Management, Greener AI
1 Introduction
The growth of convenience packaging has increased waste generation [1], underscoring the need for efficient sorting. Global waste is projected to grow from 2.1 to 3.8 billion tons by 2050 [2], an increase that compounds the financial, environmental and planetary burdens. For instance, a study on Chilean municipal solid waste (MSW) found the cost of the unsorted waste to be 297.66 euro per ton [3]. Furthermore, waste contamination poses a significant challenge to implement a circular economy, particularly given that the US has seen its recycling rate stagnate at around 35% for over a decade [4].
To address this challenge, traditional waste management has begun incorporating technology. Over the past decades, various machine learning (ML) models such as linear regression (LR), support vector machine (SVM), and random forest (RF) have evolved for predicting inbound contamination rates [5]. Simultaneously, IoT devices integrated with bins, vehicles, and recycling facilities are aiding waste sorting and data collection, leveraging GPS for route tracking and temperature sensors for fire protection. The recent shift from traditional ML to Deep Learning (DL) has delivered substantial improvements through better computation power and advanced algorithms. However, research remains uneven across sectors, often relying on simplified or artificial data [6].
The economic viability of these systems is a crucial factor, as research by Liu et al. demonstrated that computer vision-enabled systems (CVAS) become cost-effective when labor costs are high, while conventional sorting (CS) is preferable when machinery or maintenance costs are higher; their comprehensive cost model included labor, training, machinery, maintenance, and net present values of investments [7]. Special DL architectures, particular object detection models, have shown promising results in sorting applications. For example, YOLOv5 models equipped with webcam and robotic arms have demonstrated waste sorting capabilities with 93.3% accuracy [8]. More recently, a YOLOv8 model embedded with a Raspberry Pi achieved 98% accuracy in complex tasks of real-time intelligent garbage monitoring and collection systems [9] showcasing practical, low-cost solutions for smart waste management in urban environments.
In the realm of classification, high accuracy has been achieved with CNN architectures Ahmad et al. utilized ResNet-based CNN to automatically class 12 waste types, achieving 98.16% [10]. Tran et al. achieved 96% accuracy using ResNet-50 of organic and inorganic waste classification with raspberry pie 4 to direct sorting [11]; and a comparison by Soni et al. study, found MobileNet despite achieving 80% accuracy, offered a superior accuracy and lower computation cost making it highly suitable for scalable real-world applications [12].
[…]
2 Materials and Methods
The section discusses the dataset used in this study and the methods used to benchmark train the models.
2.1 Dataset and Preprocessing
In this study, we used our garbage dataset [16] focusing on a subset of seven categories deemed critical for recycling efficiency: biological, cardboard, glass, metal, paper, plastic, and trash. These images were collected from the internet, DWaste platform, and community submissions. All images were annotated with category labels and bounding boxes using Annotate Lab [17]. The final processed dataset consists of 11,163 images and 19,700 bounding box instances shown below Figure 1.
Figure 1: Dataset class distribution
The above images show a non-uniform distribution characteristic of real world municipal solid waste streams. For classification models, class imbalance was addressed using undersampling technique, where images were selectively removed from oversampled classes [18]. Conversely, for object detection models, the imbalance was addressed by applying computed class weights during the training phase [19], which up-weighted the loss contribution from underrepresented classes. The finalized dataset was then partitioned using an 80/20 split for the training and validation sets. The sample annotated waste image of the above classes is shown in Figure 2.
Figure 2: Sample annotated images from our dataset
2.2 Model Training
We evaluated both classification (EfficientNetV2S/M, MobileNet, ResNet50/101) and object detection (YOLOv8n, YOLOv11n) architectures using a transfer learning approach. All classification models were initialized with weight pre-trained on the ImageNet dataset. All models were trained for 20 epochs using NVIDIA Tesla T4x2 GPU in Kaggle. The performance of each model was benchmarked using standard performance metrics including Accuracy, Precision, Recall, F1-score, and Mean Average Precision (mAP). With our focus on sustainable AI deployment, we conducted detailed analysis of VRAM usage, model size, and carbon emissions across various phases of the workflow. Carbon emissions were monitored using CodeCarbon library [20]. The resulting full-precision models were further optimized via quantization, a technique known to achieve up to 95% reduction in the number of parameters and model size [21], thereby significantly lowering energy use. Figure 3 depicts the overall process, while the corresponding experiments are available on Github [^1] https://github.com/sumn2u/AIMS-2025.
Figure 3: End-to-end machine learning pipeline
3 Results and Discussion
Our experiment revealed a distinct tradeoff between model accuracy, size, and carbon efficiency as shown in Table 1.
Table 1: Experimental Results (Performance metrics and Model Sizes).
Model (Classification) Acc (%) P R F1 Size (MB) Q-Size (MB)
MobileNet 67.50 0.67 0.68 0.67 14.7 3.5
EffNetV2M 94.70 0.94 0.95 0.95 216.0 56.4
EffNetV2S 96.00 0.96 0.96 0.96 84.3 22.1
ResNet101 92.10 0.91 0.93 0.92 174.6 43.6
ResNet50 91.40 0.90 0.92 0.91 97.9 24.2
Metrics: Acc = Accuracy, P = Precision, R = Recall, F1 = F1 Score.
Model (Detection) Acc (%) P R F1 mAP Size (MB) Q-Size (MB)
YOLOv8n 75.50 0.78 0.65 0.75 0.76 6.5 3.1
YOLOv11n 77.10 0.77 0.69 0.77 0.77 5.4 2.8
Metrics: mAP = mean Average Precision. Q-Size = Quantized Size. Best results are in bold.
The high-performance classification models EfficientNetV2M and EfficientNetV2S showed highest accuracy ($\approx 95-96\%$) but was constrained by larger model sizes (216MB and 84.3MB respectively) and consequently resulted in higher training and deployment emissions shown in Figure 4. Similarly, ResNet101 and ResNet50 delivered strong accuracy (91-92%) but were also penalized by substantial size and higher emissions. In contrast, MobileNet prioritized resource efficiency, exhibiting the smallest initial size (14.7 MB, reduced to 3.5 MB after quantization) with minimal energy usage and lowest accuracy ($\approx 67\%$). The lightweight object detection models, YOLOv8n and YOLOv11n demonstrated the most balanced trade-off, achieving accuracy $75\text{-}77\%$ accuracy with impressive quantized sizes ($\approx 3$MB).
Figure 4: Carbon emission by stage for each model
While all classification models, including MobileNet, exhibited low VRAM usage during inference time as shown in Figure 5 and Figure 6. In contrast, the YOLO models consumed slightly more VRAM initially as shown in Figure 7 but achieved faster inference time, which further improved significantly following quantization.
Figure 5: Classification models VRAM usage under inference
Figure 6: Object detection models VRAM usage under inference
Figure 7: Object detection models VRAM usage after quantization under inference
Notably, YOLOv11n emerged as the optimal architecture for edge deployment, attaining the best mean Average Precision (mAP = 0.77) with the smallest quantized footprint (2.8 MB) and the fastest speed, confirming its efficacy for resource-constrained applications with minimal carbon emissions.
This optimized model has been successfully deployed to both the DWaste mobile application and a dedicated edge device for a real time waste object. An example of waste detection using both the app and edge device is shown in Figure 8.
Figure 8: Real-world detection on edge and mobile device
4 Conclusion
This study successfully quantified the inherent trade-off between model accuracy and energy efficiency for deep learning-based waste sorting systems. Our results clearly demonstrate that while larger architectures (EfficientNetV2S/M, ResNet101/50) achieve superior classification accuracy, they demand greater computational resources and incur significantly higher carbon emissions. Conversely, lightweight object detection models, specifically YOLOv11n and YOLOv8n, strike a crucial balance, offering strong real-time performance (mAP 77%) with minimal resource overhead. Furthermore, we confirmed that the application of model quantization dramatically aids deployment, consistently and substantially reducing VRAM usage and model size, thus directly lowering the energy required for inference on edge and mobile devices. Given these findings, the YOLOv11n model, which achieved the best combination of accuracy, inference speed, and minimal resource usage (smallest quantized footprint), was selected and successfully embedded into the DWaste mobile app and a dedicated edge device for practical waste sorting implementation. Future work should focus on refining lightweight model accuracy with an expanded dataset and conducting a longitudinal study to evaluate the long-term effectiveness and economic impact of the deployed app and the edge system.